OpenStudy (anonymous):

Use the graph to find the solution to the system of equations. Enter your answer as an ordered pair. y=x+3 y=3x+1

OpenStudy (anonymous):

Where is the graph ?

OpenStudy (anonymous):

You are receiving this error for one of the following reasons: 1. This activity has been inactive for too long and has shut down. Solution: Try accessing the activity again from the page where you originally launched it. If that doesn't work, go back to your virtual school web site and log in again. 2. You are accessing your course material through a bookmark. Solution: Close down your current browsers, and go back to your virtual school web site and log in again. 3. Another person logged in with your username and password. Solution: Close down your current browsers, and go back to your virtual school web site and log in again. 4. You are using the AOL browser to access your course material. Solution: The AOL browser doesn't always perform satisfactorily. AOL users should use AOL to connect to the Internet and then use Internet Explorer to access your online learning course material. Follow these steps: 1. Connect to the internet using AOL. 2. Minimize the AOL browser. 3. Open Internet Explorer and access your course material. If you continue to have problems or have any other questions please contact Apex Learning Support. By email: support@apexlearning.com By phone: 1-800-453-1454 Cookie Value = 4043 Server Value = 4036

OpenStudy (anonymous):

oh... ok well I cant post the graph. but I have the equations.

OpenStudy (anonymous):

x=1; y=4

OpenStudy (anonymous):

thank u!

OpenStudy (anonymous):

y=x+3,x=0,y=0+3=3 when y=0,x=-3 y=3x+1 when x=0,y=0+1=1 when x=-1,y=-3+1=-2 

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

solution written on the graph is wrongly written as (1,3) but correct solution is (1,4)